merisaa is 2 years old her father is 16 times her age & how many years will he be twice is age
Answer by stanbon(75887)

merisaa is 2 years old her father is 16 times her age & how many years will he be twice her age
----
Now DATA:
Marissa is 2 yrs old.
Her father is 16*2 = 32 yrs old.
--------------------------
Equation:
32+x = 2(2+x)
32+x = 4+2x
x = 28
----
In 28 years the father will be 60 and Marissa will be 30.
==============================================================
Cheers,
Stan H.